### v1

# Typical infinite mixing automorphisms are rank one

Typical infinite mixing automorphisms are rank one
Valery V. Ryzhikov
Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric have rank one. Using Sidon’s constructions, we showed that generic infinite mixing automorphisms also have rank one.
1 Introduction
Alpern and Tikhonov [6] introduced a complete metric on the space of mixing actions. Bashtanov [1] proved that the class of rank-one actions is generic in the space ${\bf Mix}$ of invertible mixing transformations with invariant probability measure. We will establish a similar theorem for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$ of transformations of $T$ with sigma-finite invariant measure. We will prove that the conjugacy class of the Sidon transformation of rank one (see [4], [5]) is dense in ${\bf Mix_{\infty}}$. Since the class of rank-one transformations is a $G_{\delta}$-set, taking into account the density of the latter, we obtain the genericity of this class. A consequence of this typicality, given the result [2], is the typicality of the minimality of the centralizer (the transformation $T$ commutes only with its powers).
The group ${\bf Aut_{\infty}}$ of all atomorphisms of the space $(X,\cal B,\mu)$ isomorphic to the line ${\mathbf{R}}$ with the standard Lebesgue measure is equipped with the Halmos metric $\rho$:
$\rho(S,T)=\sum_{i}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right),$
where $\{A_{i}\}$ is some fixed family dense in the family of all measurable sets of finite measure. The space $({\bf Aut_{\infty}},\rho)$ is a complete separable space. It is easy to show that in $({\bf Aut_{\infty}},\rho)$ the so-called mixing transformations form a set of the first category (countable intersection of nowhere dense sets). If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$). The question arises: which invariants, i.e. properties preserved under conjugation in $\bf Aut_{\infty}$) are typical? A property is called typical if a dense $G_{\delta}$-set of transformations has this property.
For example, given the known methods, it can be shown that in ${\bf Mix}_{\infty}$ the spectrum of all symmetric powers of the transformation is typically single. It easily follows from the results of [5] that a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
Spaces $\bf{\bf Mix}$, ${\bf Mix}_{\infty}$. Following [6], we define the metric $d_{w}$ on ${\bf Aut}(\mu)$:
$d_{w}(S,T)=\sum_{i,j}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
On the set of all mixing automorphisms Mix is defined the metric $r$:
$r(S,T)=\rho(S,T)+\sup_{n>0}d_{w}(S^{n},T^{n}).$
In [6] it is proved that the metric space $({\bf Mix},r)$ is complete and separable.
Consider the sigma-finite measure $\mu_{\infty}$ as an invariant and, without changing the formulas, introduce the same metric $r$ on the space of mixing infinite automorphisms:
${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty}),\,T^{n}\to_{w}0\},$
where $\{A_{i}\}$ is a fixed sequence of sets, dense in the family of all sets of finite measure.
The sequence $T_{k}$, fundamental in $({\bf Mix_{\infty}},r)$, converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$. It is easy to show that in this case $T_{k}\to T$ and in the metric $\rho$.
2 Density of the conjugacy class of a Sidon automorphism
Let us recall the necessary definitions.
Construction of rank one. Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of superstructures)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})).$
At step $j=1$, by definition, a set of half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ is given. Let at step $j$ a system of non-intersecting half-intervals
$E_{j},SE_{j},S^{2}E_{j},\dots,S^{h_{j}-1}E_{j},$
be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is a parallel translation. Such a set of half-intervals is called the tower of step $j$, their union is denoted by $X_{j}$ and is also called a tower.
We represent $E_{j}$ as a disjoint union of $r_{j}$ half-intervals
$E_{j}^{1},E_{j}^{2}E_{j}^{3},\dots E_{j}^{r_{j}}$
of the same measure (length). For each $i=1,2,\dots,r_{j}$, we consider the so-called column
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i}.$
We denote the union of these half-intervals by $X_{i,j}$, which we will also call a column.
To each column $X_{i,j}$ we add $s_{j}(i)$ disjoint half-intervals of the same measure as $E_{j}^{i}$, obtaining a set
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i},S^{h_{j}}E_{j}^{i},S^{h_{j}+1}E_{j}^{i},\dots,S^{h_{j}+s_{j}(i)-1}E_{j}^{i}$
(all these sets are disjoint). Denoting $E_{j+1}=E^{1}_{j}$, for $i<r_{j}$ we set
$S^{h_{j}+s_{j}(i)}E_{j}^{i}=E_{j}^{i+1}.$
The set of superstructured columns is from now on considered as a tower of stage $j+1$, consisting of half-intervals
$E_{j+1},SE_{j+1},S^{2}E_{j+1},\dots,S^{h_{j+1}-1}E_{j+1},$
where
$h_{j+1}+1=(h_{j}+1)r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
The partial definition of the transformation $S$ at stage $j$ is preserved at all subsequent stages. As a result, on the space $X=\cup_{j}X_{j}$ there is defined an invertible transformation $S:X\to X$ that preserves the standard Lebesgue measure on $X$.
Let the construction $S$ of rank one have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$. Such a construction is called Sidon. The measure of the space $X$ in this case is infinite.
Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
Lemma 2. Let $T$ be a mixing transformation, $A$ a set of finite measure. The phase space for every natural $H$ can be represented as
$X=\bigsqcup_{h>H}\bigsqcup_{i=0}^{h}T^{i}B_{h},\ \ \ T^{h}B_{h}\cap A=\phi.$
The proof of the lemma is simple, we leave it as an exercise.
The proof of the theorem is brief. Let us consider a part of the Kakutani tower:
$U_{N}=\bigsqcup_{H<h<N}\bigsqcup_{i=0}^{h}T^{i}B_{h},$
which almost absorbed the sets $A_{1},\dots,A_{k}$, appearing in the definition of the metric $r$. For this, we choose a large $N$.
We take a very thin tower $X_{j}$, appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in such a way that the following conditions are satisfied:
(1) the conjugate transformation $\tilde{S}$ coincides with $T$ on $U_{N}$ minus the upper floors of $T^{h}B_{h}$ and minus a set of arbitrarily small measure;
(2) after the moving floor of the wound tower leaves the top of $U_{N}$ for the $k$-th time, it must travel outside $U_{N}$ for a time $s_{k}>10s_{k-1}$ ($s_{1}>10N$). This method can be called the Sidon winding, which will ensure that the intersection measure is small:
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap U_{N})\ <\ N\mu(B_{j}),\ \ m<h_{j}.$
For $m\geq h_{j}$, the inequality
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap\tilde{U}_{N})<\mu(\tilde{U}_{N})/r_{j}$
is satisfied by the definition of the Sidon construction.
If $A=A_{1}\cup\dots\cup A_{k}\subset\tilde{U}_{N}$, then
$\sum_{i}^{k}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right)=0$
and
$\sum_{i,j=1}^{k}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|=0.$
In reality, when choosing a large $N$ and a suitable winding, instead of equalities we get inequalities with a very small right-hand side.
So for $\varepsilon>0$ there are always $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence are close in the Halmos metric, but their degrees are uniformly close with respect to the $d_{w}$ metric. Thus we obtain $r(R^{-1}SR,T)<\varepsilon$ for a suitable $R$.
The set of rank-one transformations is a $G_{\delta}$ set (we use the proof from [1] without any changes). We have found that it is dense. Thus, the typicality of rank-one is proved.
In [3] we used a similar trick, but the windings were Ornstein and the spacers for flows and automorphisms of rank one were Ornstein as well. In the case of infinite measure, a pleasant opportunity arises to use large Sidon constructions, which allows us to use a simpler and more constructive method. For other results on Sidon transformations, see [4].
3 Типичность ранга один в ${\bf Mix_{\infty}}$
Альперн и Тихонов [6] ввели полную метрику на пространстве перемешивающих действий. Баштанов [1] доказал, что класс действий ранга один типичен в пространстве ${\bf Mix}$братимых перемешивающих преобразований с инвариантной вероятностной мерой. Мы установим аналогичную теорему для пространства ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$реобразований $T$игма-конечной инвариантной мерой. Докажем, что класс сопряженности сидоновского преобразования ранга один (см. [4], [5]) плотен в ${\bf Mix_{\infty}}$. Так как класс преобразований ранга один является $G_{\delta}$-множеством, с учетом плотности последнего получаем типичность этого класса. Следствием этой типичности с учетом результата [2] является типичность минимальности централизатора (преобразование $T$оммутирует только со своими степенями).
Группа ${\bf Aut_{\infty}}$сех атоморфизмов пространства $(X,\cal B,\mu)$, изоморфного прямой ${\mathbf{R}}$о стандартной мерой Лебега оснащается метрикой Халмоша $\rho$:
$\rho(S,T)=\sum_{i}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right),$
где $\{A_{i}\}$ – некоторое фиксированное семейство, плотное в семействе всех измеримых множеств конечной меры. Пространство $({\bf Aut_{\infty}},\rho)$вляется полным сепарабельным пространством. Несложно показать, что в $({\bf Aut_{\infty}},\rho)$ак называемые перемешивающие преобразования образуют множство первой категории (счетное пересечение нигде не плотных множеств). Если же усилить метрику Халмоша до метрики Альперна-Тихонова, получим полное метрическое пространство ${\bf Mix}_{\infty}$). Возникает вопрос: какие инварианты, т.е. свойства, сохранящиеся при сопряжении в $\bf Aut_{\infty}$) являются типичными? Свойство называется типичным, если плотное $G_{\delta}$-множество преобразований обладают этим свойством.
Например, с учетом известных методов можно показать, что в ${\bf Mix}_{\infty}$ипична однократность спектра всех симметрических степеней преобрзования. Из результатов [5] легко вытекает, что типичное в ${\bf Mix}_{\infty}$реобразование не изоморфно своему обратному.
Пространства $\bf{\bf Mix}$, ${\bf Mix}_{\infty}$. Cледуя [6], определим на ${\bf Aut}(\mu)$етрику $d_{w}$:
$d_{w}(S,T)=\sum_{i,j}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
На множестве всех перемешиващих автоморфизмов Mix задаем метрику $r$:
$r(S,T)=\rho(S,T)+\sup_{n>0}d_{w}(S^{n},T^{n}).$
В [6] доказано, что метрическое пространство $({\bf Mix},r)$вляется полным и сепарабельным.
Рассмотрим в качестве инвариантной сигма-конечную меру $\mu_{\infty}$, не меняя формул, введем такую же метрику $r$а пространстве перемешивающих бесконечных автоморфизмов:
${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty}),\,T^{n}\to_{w}0\},$
где $\{A_{i}\}$ – фиксированная последовательность множеств, плотная в семействе всех множеств конечной меры.
Фундаментальная в $({\bf Mix_{\infty}},r)$оследовательность $T_{k}$ходится к некоторому $T$илу полноты метрики Халмоша в пространстве ${\bf Aut}(\mu_{\infty})$. Несложно показать, что в этом случае $T_{k}\to T$етрике $\rho$.
Плотность класса сопряженности сидоновского автоморфизма.
Конструкция ранга один.
Фиксируем натуральное число $h_{1}$, последовательность $r_{j}\to\infty$ ( параметр $r_{j}$ – число колонн, на которые разрезается башня этапа $j$) и последовательность целочисленных векторов (параметров надстроек)
$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})).$
На шаге $j=1$о определению задан набор полуинтервалов $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$. Пусть на шаге $j$пределена система непересекающихся полуинтервалов
$E_{j},SE_{j},S^{2}E_{j},\dots,S^{h_{j}-1}E_{j},$
причем на полуинтервалах $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ребразование $S$вляется параллельным переносом. Такой набор полуинтервалов называется башней этапа $j$, их объединение обозначается через $X_{j}$оже называется башней.
Представим $E_{j}$ак дизъюнктное объединение $r_{j}$олуинтервалов
$E_{j}^{1},E_{j}^{2}E_{j}^{3},\dots E_{j}^{r_{j}}$
одинаковой меры (длины). Для каждого $i=1,2,\dots,r_{j}$ассмотрим так называемую колонну
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i}.$
Объдинение этих полуинтервалов обозначим через $X_{i,j}$, его тоже будем называть колонной.
К каждой колонне $X_{i,j}$обавим $s_{j}(i)$епересекающихся полуинтервалов той же меры, что у $E_{j}^{i}$, получая набор
$E_{j}^{i},SE_{j}^{i},S^{2}E_{j}^{i},\dots,S^{h_{j}-1}E_{j}^{i},S^{h_{j}}E_{j}^{i},S^{h_{j}+1}E_{j}^{i},\dots,S^{h_{j}+s_{j}(i)-1}E_{j}^{i}$
(все эти множества не пересекаются). Обозначив $E_{j+1}=E^{1}_{j}$, для $i<r_{j}$оложим
$S^{h_{j}+s_{j}(i)}E_{j}^{i}=E_{j}^{i+1}.$
Набор надстроеных колонн с этого момента рассматривается как башня этапа $j+1$, состоящая из полуинтервалов
$E_{j+1},SE_{j+1},S^{2}E_{j+1},\dots,S^{h_{j+1}-1}E_{j+1},$
где
$h_{j+1}+1=(h_{j}+1)r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
Частичное определение преобразования $S$а этапе $j$охраняется на всех следующих этапах. В итоге на пространстве $X=\cup_{j}X_{j}$пределено обратимое преобразование $S:X\to X$, сохраняющее стандартную меру Лебега на $X$.
Пусть конструкция $S$анга один обладает следующим свойством: пересечение $X_{j}\cap S^{m}X_{j}$ри $h_{j}<m\leq h_{j+1}$ожет содержаться только в одной из колонн $X_{i,j}$ашни $X_{j}$. Такая конструкция называется сидоновской. Мера пространства $X$том случае бесконечна.
Теорема 1. Пусть $T$ – перемешивающее преобразование, $S$ – сидоновское преобразование ранга один. Для всякого $\varepsilon>0$айдется такое сопряжение $R$, что $r(R^{-1}SR,T)<\varepsilon$.
Лемма 2. Пусть $T$ – перемешивающее преобразование, $A$ – множество конечной меры. Фазовое пространство для всякого натурального $H$ожно представить в виде
$X=\bigsqcup_{h>H}\bigsqcup_{i=0}^{h}T^{i}B_{h},\ \ \ T^{h}B_{h}\cap A=\phi.$
Доказательство леммы несложное, оставляем его в качестве упражнения.
Доказательство теоремы кратко. Рассмотрим часть башни Какутани:
$U_{N}=\bigsqcup_{H<h<N}\bigsqcup_{i=0}^{h}T^{i}B_{h},$
которая почти поглотила множества $A_{1},\dots,A_{k}$, фигурирующие в определении метрики $r$. Для этого выбираем большое $N$.
Возьмем очень тонкую башню $X_{j}$, фигурирующую в построении сидоновского преобразования $S$, намотаем на $U_{N}$оответствующую сопряжению $\tilde{S}=R^{-1}SR$ашню $\tilde{X}_{j}$аким образом, чтобы выполнялись условия:
(1) сопряженное преобразование $\tilde{S}$овпадало с $T$а $U_{N}$инус верхние этажи $T^{h}B_{h}$инус множество сколь угодно малой меры;
(2) после того как движущийся этаж намотанной башни покидает в $k$-тый раз верх $U_{N}$, он должен путешествовать вне $U_{N}$ремя $s_{k}>10s_{k-1}$ ($s_{1}>10N$). Этот способ можно назавать сидоновской намоткой, которая обеспечит нам малость меры пересечения:
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap U_{N})\ <\ N\mu(B_{j}),\ \ m<h_{j}.$
При $m\geq h_{j}$ыполняется неравенство
$\mu(\tilde{S}^{m}\tilde{U}_{N}\cap\tilde{U}_{N})<\mu(\tilde{U}_{N})/r_{j}$
в силу определения сидоновской конструции.
Если бы выполнялось $A=A_{1}\cup\dots\cup A_{k}\subset\tilde{U}_{N}$, то
$\sum_{i}^{k}2^{-i}\left(\mu(SA_{i}\Delta TA_{i})+\mu(S^{-1}A_{i}\Delta T^{-1}A_{i})\right)=0$
и
$\sum_{i,j=1}^{k}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|=0.$
В действительности при выборе большого $N$одходящей обмотке вместо равенств получим неравенства с очень малой правой частью.
Итак для $\varepsilon>0$сегда найдутся $j$, $N$ $R$акие, что на пересечении $\tilde{X}_{j}\cap U_{N}$реобразования $\tilde{S}$ $T$очти совпадают, значит близки в метрике Халмоша, но их степени равномерно близки относительно метрики $d_{w}$. Тем самым мы получим $r(R^{-1}SR,T)<\varepsilon$ля подходящего $R$.
Множество преобразований ранга один является $G_{\delta}$ножеством (используем без особых изменений доказательство из [1]). Мы получили, что оно плотное. Таким образом, типичность ранга один доказана.
Замечание. В [3] мы использовали подобный прием, но обмотки были орнстейновскими и надстройки для потоков и автоморфизов ранга один были орнстейновскими. В случае бесконечной меры возникает приятная возможность использовать большие сидоновкие надстройки, что позволяет использовать более простой и конструктивный метод. Другие результаты о сидоновских преобразованиях см. [4].
Список литературы
  • [1] А.И. Баштанов, Типичное перемешивание имеет ранг 1, Матем. заметки, 93:2 (2013), 163-171; A.I. Bashtanov, Generic Mixing Transformations Are Rank 1, Math. Notes, 93:2 (2013), 209-216
  • [2] В.В. Рыжиков, Ж.-П. Тувено, О централизаторе бесконечного перемешивающего преобразования ранга один, Функц. анализ и его прил., 49:3 (2015), 88-91; V.V. Ryzhikov, J.-P. Thouvenot, On the Centralizer of an Infinite Mixing Rank-One Transformation, Funct. Anal. Appl., 49:3 (2015), 230-233
  • [3] В.В. Рыжиков, О сохраняющих меру преобразованиях ранга один, Тр. ММО, 81, № 2, МЦНМО, М., 2020, 281-318; V.V. Ryzhikov, Measure-preserving rank one transformations, Trans. Moscow Math. Soc., 81:2 (2020), 229-259
  • [4] В.В. Рыжиков, Полиномиальная жесткость и спектр сидоновских автоморфизмов. Матем. сб., 215:7 (2024), 138-152; V.V. Ryzhikov, Polynomial rigidity and spectrum of Sidon automorphisms, Mat. Sb., 215:7 (2024), 138-152
  • [5] V.V. Ryzhikov, Asymmetric mixing Poisson suspensions, Arxiv 2406.14390
  • [6] С.В. Тихонов, Полная метрика на множестве перемешивающих преобразований, Матем. сб., 198:4 (2007), 135-158; S.V. Tikhonov, A complete metric in the set of mixing transformations, Sb. Math., 198:4 (2007), 575-596

### v2

@@
-Typical infinite mixing automorphisms are rank one
-Typical infinite mixing automorphisms are rank one
+Typical infinite mixing automorphisms are rank-one
+Typical infinite mixing automorphisms are rank-one
 Valery V. Ryzhikov
-Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric have rank one.
-Using Sidon’s constructions, we showed that generic infinite mixing automorphisms also have rank one.
+Bashtanov proved that generic mixing automorphisms of probability space with respect to the Alpern-Tikhonov metric had rank one.
+Using Sidon’s constructions, we show that generic infinite mixing automorphisms also are rank-one.
 1 Introduction
-Alpern and Tikhonov [6] introduced a complete metric on the space of mixing actions.
+Alpern and Tikhonov [1],[7] introduced a complete metric on the space of mixing actions.
 Bashtanov [2] proved that the class of rank-one actions is generic in the space ${\bf Mix}$ of invertible mixing transformations with invariant probability measure.
-We will establish a similar theorem for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to 0\}$ of transformations of $T$ with sigma-finite invariant measure.
-We will prove that the conjugacy class of the Sidon transformation of rank one (see [4], [5]) is dense in ${\bf Mix_{\infty}}$.
+We will establish the same for the space ${\bf Mix_{\infty}}=\{T:\,T^{n}\to_{w}0\}$ of transformations of $T$ with sigma-finite invariant measure.
+We prove that the conjugacy class of the Sidon transformation of rank one (see [5], [6]) is dense in ${\bf Mix_{\infty}}$.
 Since the class of rank-one transformations is a $G_{\delta}$-set, taking into account the density of the latter, we obtain the genericity of this class.
-A consequence of this typicality, given the result [2], is the typicality of the minimality of the centralizer (the transformation $T$ commutes only with its powers).
+A consequence of this and the result of [3] is the typicalness of the minimality of the centralizer (the tipical transformation $T$ commutes only with its powers).
 The group ${\bf Aut_{\infty}}$ of all atomorphisms of the space $(X,\cal B,\mu)$ isomorphic to the line ${\mathbf{R}}$ with the standard Lebesgue measure is equipped with the Halmos metric $\rho$:
@@
 where $\{A_{i}\}$ is some fixed family dense in the family of all measurable sets of finite measure.
-The space $({\bf Aut_{\infty}},\rho)$ is a complete separable space.
+The space $({\bf Aut_{\infty}},\rho)$ is a complete separable metric space.
 It is easy to show that in $({\bf Aut_{\infty}},\rho)$ the so-called mixing transformations form a set of the first category (countable intersection of nowhere dense sets).
-If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$).
-The question arises: which invariants, i.e.
+If we strengthen the Halmos metric to the Alpern-Tikhonov metric, we obtain a complete metric space ${\bf Mix}_{\infty}$.
+Which invariants (i.e.
 properties preserved under conjugation in $\bf Aut_{\infty}$) are typical?
 A property is called typical if a dense $G_{\delta}$-set of transformations has this property.
-For example, given the known methods, it can be shown that in ${\bf Mix}_{\infty}$ the spectrum of all symmetric powers of the transformation is typically single.
-It easily follows from the results of [5] that a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
 Spaces $\bf{\bf Mix}$, ${\bf Mix}_{\infty}$. Following [7], we define the metric $d_{w}$ on ${\bf Aut}(\mu)$:
-$d_{w}(S,T)=\sum_{i,j}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
+$d_{w}(S,T)=\sum_{i,j=1}^{\infty}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|.$
 On the set of all mixing automorphisms Mix is defined the metric $r$:
@@
 Consider the sigma-finite measure $\mu_{\infty}$ as an invariant and, without changing the formulas, introduce the same metric $r$ on the space of mixing infinite automorphisms:
-${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty}),\,T^{n}\to_{w}0\},$
+${\bf Mix}_{\infty}=\{T\in{\bf Aut}(\mu_{\infty})\,:\,T^{n}\to_{w}0\},$
 where $\{A_{i}\}$ is a fixed sequence of sets, dense in the family of all sets of finite measure.
-The sequence $T_{k}$, fundamental in $({\bf Mix_{\infty}},r)$, converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$.
-It is easy to show that in this case $T_{k}\to T$ and in the metric $\rho$.
-2 Density of the conjugacy class of a Sidon automorphism
-Let us recall the necessary definitions.
+Fundamental in $({\bf Mix_{\infty}},r)$ a sequence $T_{k}$ converges to some $T$ by virtue of the completeness of the Halmos metric in the space ${\bf Aut}(\mu_{\infty})$.
+It is easy to show that this $T_{k}\to T$ converges to $T$ in the metric $r$ as well.
+From the known methods, it can be shown that
+for generic transformations in ${\bf Mix}_{\infty}$ that spectra of their symmetric powers are simple.
+It follows from the results of [6] that
+a typical transformation in ${\bf Mix}_{\infty}$ is not isomorphic to its inverse.
+2 Density of the conjugacy classes of the Sidon automorphisms
+Let us recall necessary definitions.
 Construction of rank one.
-Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of superstructures)
-$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})).$
-At step $j=1$, by definition, a set of half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ is given. Let at step $j$ a system of non-intersecting half-intervals
+Fix a natural number $h_{1}$, a sequence $r_{j}\to\infty$ (parameter $r_{j}$ is the number of columns into which the tower of stage $j$ is cut) and a sequence of integer vectors (parameters of spacers)
+$\bar{s}_{j}=(s_{j}(1),s_{j}(2),\dots,s_{j}(r_{j}-1),s_{j}(r_{j})),\ s_{(}i)\geq 0.$
+At step $j=1$ a set of disjoint half-intervals $E_{1},SE_{1},S^{2}E_{1},\dots,S^{h_{1}-1}E_{1}$ of the same measure is given.
+Let at step $j$ a system of non-intersecting half-intervals
 $E_{j},SE_{j},S^{2}E_{j},\dots,S^{h_{j}-1}E_{j},$
-be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is a parallel translation.
+be defined, and on the half-intervals $E_{j},SE_{j},\dots,S^{h_{j}-2}E_{j}$ the transformation $S$ is the parallel translation.
 Such a set of half-intervals is called the tower of step $j$, their union is denoted by $X_{j}$ and is also called a tower.
@@
 where
-$h_{j+1}+1=(h_{j}+1)r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
+$h_{j+1}=h_{j}r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
 The partial definition of the transformation $S$ at stage $j$ is preserved at all subsequent stages.
 As a result, on the space $X=\cup_{j}X_{j}$ there is defined an invertible transformation $S:X\to X$ that preserves the standard Lebesgue measure on $X$.
-Let the construction $S$ of rank one have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$.
+Sidon constructions.
+Let a rank one construction $S$ have the following property: the intersection $X_{j}\cap S^{m}X_{j}$ for $h_{j}<m\leq h_{j+1}$ can be contained only in one of the columns $X_{i,j}$ of the tower $X_{j}$.
 Such a construction is called Sidon.
-The measure of the space $X$ in this case is infinite.
-Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
+The measure of its space $X$ is infinite.
+If $r_{j}\to\infty$, then Sidon $S$ is mixing.
+Theorem 1. Let $T$ be a mixing transformation, $S$ a Sidon mixing transformation of rank one. For every $\varepsilon>0$ there exists a conjugation $R$ such that $r(R^{-1}SR,T)<\varepsilon$.
 Lemma 2. Let $T$ be a mixing transformation, $A$ a set of finite measure. The phase space for every natural $H$ can be represented as
 $X=\bigsqcup_{h>H}\bigsqcup_{i=0}^{h}T^{i}B_{h},\ \ \ T^{h}B_{h}\cap A=\phi.$
-The proof of the lemma is simple, we leave it as an exercise.
-The proof of the theorem is brief. Let us consider a part of the Kakutani tower:
+The proof of the lemma is simple, let it be an exercise.
+The proof of the theorem (in brief).
+Let us consider a part of the Kakutani tower:
 $U_{N}=\bigsqcup_{H<h<N}\bigsqcup_{i=0}^{h}T^{i}B_{h},$
 which almost absorbed the sets $A_{1},\dots,A_{k}$, appearing in the definition of the metric $r$. For this, we choose a large $N$.
-We take a very thin tower $X_{j}$, appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in such a way that the following conditions are satisfied:
+We take a very thin tower $X_{j}$ (so very large $j$), appearing in the construction of the Sidon transformation $S$, and wind on $U_{N}$ the tower $\tilde{X}_{j}$ corresponding to the conjugation $\tilde{S}=R^{-1}SR$ in a way that the following conditions are satisfied:
 (1) the conjugate transformation $\tilde{S}$ coincides with $T$ on $U_{N}$ minus the upper floors of $T^{h}B_{h}$ and minus a set of arbitrarily small measure;
 (2) after the moving floor of the wound tower leaves the top of $U_{N}$ for the $k$-th time, it must travel outside $U_{N}$ for a time $s_{k}>10s_{k-1}$ ($s_{1}>10N$).
-This method can be called the Sidon winding, which will ensure that the intersection measure is small:
+This Sidon winding will ensure that the intersection measure is small:
 $\mu(\tilde{S}^{m}\tilde{U}_{N}\cap U_{N})\ <\ N\mu(B_{j}),\ \ m<h_{j}.$
@@
 $\sum_{i,j=1}^{k}2^{-i-j}\left|\mu(SA_{i}\cap A_{j})-\mu(TA_{i}\cap A_{j})\right|=0.$
-In reality, when choosing a large $N$ and a suitable winding, instead of equalities we get inequalities with a very small right-hand side.
-So for $\varepsilon>0$ there are always $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence are close in the Halmos metric, but their degrees are uniformly close with respect to the $d_{w}$ metric.
+In reality, for a large $N$ and a suitable Sidon winding, instead of equalities we get inequalities with a very small right-hand side.
+So for $\varepsilon>0$ there are $j$, $N$ and $R$ such that at the intersection $\tilde{X}_{j}\cap U_{N}$ the transformations $\tilde{S}$ and $T$ almost coincide, hence, $\tilde{S}$ and $T$ are close in the Halmos metric, but their powers are uniformly close with respect to the $d_{w}$ metric.
 Thus we obtain $r(R^{-1}SR,T)<\varepsilon$ for a suitable $R$.
-The set of rank-one transformations is a $G_{\delta}$ set (we use the proof from [1] without any changes). We have found that it is dense. Thus, the typicality of rank-one is proved.
-In [3] we used a similar trick, but the windings were Ornstein and the spacers for flows and automorphisms of rank one were Ornstein as well.
-In the case of infinite measure, a pleasant opportunity arises to use large Sidon constructions, which allows us to use a simpler and more constructive method.
-For other results on Sidon transformations, see [4].
+The set of rank-one transformations is a $G_{\delta}$ set (use the proof from [2]). We have found that the class of mixing rank one transfomations is dense in $\bf Mix_{\infty}$. Thus, it is typical.
+In [4] we used a similar trick, but the windings were random Ornstein and the spacers for flows and automorphisms of rank one were random Ornstein as well.
+In the case of infinite measure, a pleasant opportunity arises to use large Sidon spacers, which gives us a simpler and constructive method. For other results on Sidon transformations, see in [5].
 3 Типичность ранга один в ${\bf Mix_{\infty}}$
-Альперн и Тихонов [6] ввели полную метрику на пространстве перемешивающих действий.
+Альперн и Тихонов [1],[7] ввели полную метрику на пространстве перемешивающих действий.
 Баштанов [2] доказал, что класс действий ранга один типичен в пространстве ${\bf Mix}$братимых перемешивающих преобразований с инвариантной вероятностной мерой.
@@
 где
-$h_{j+1}+1=(h_{j}+1)r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
+$h_{j+1}=h_{j}r_{j}+\sum_{i=1}^{r_{j}}s_{j}(i).$
 Частичное определение преобразования $S$а этапе $j$охраняется на всех следующих этапах.
@@
 Тем самым мы получим $r(R^{-1}SR,T)<\varepsilon$ля подходящего $R$.
-Множество преобразований ранга один является $G_{\delta}$ножеством (используем без особых изменений доказательство из [1]). Мы получили, что оно плотное. Таким образом, типичность ранга один доказана.
+Множество преобразований ранга один является $G_{\delta}$ножеством (используем без особых изменений доказательство из [2]). Мы показали его плотность. Таким образом, типичность ранга один доказана.
 Замечание.
-В [3] мы использовали подобный прием, но обмотки были орнстейновскими и надстройки для потоков и автоморфизов ранга один были орнстейновскими.
-В случае бесконечной меры возникает приятная возможность использовать большие сидоновкие надстройки, что позволяет использовать более простой и конструктивный метод.
+В [4] мы использовали подобный прием, но обмотки были случайными орнстейновскими и надстройки для потоков и автоморфизов ранга один также были случайными орнстейновскими.
+В случае бесконечной меры возникает приятная возможность использовать сидоновкие надстройки, что дает простой и конструктивный метод.
 Другие результаты о сидоновских преобразованиях см.
-[4].
+в [5].
 Список литературы
+  • [1] Alpern St., Conjecture: In general a mixing transformation is not two-fold mixing, Ann. Probab., 13:1 (1985), 310-313
   • [2] А.И. Баштанов, Типичное перемешивание имеет ранг 1, Матем. заметки, 93:2 (2013), 163-171; A.I. Bashtanov, Generic Mixing Transformations Are Rank 1, Math. Notes, 93:2 (2013), 209-216